OpenStudy (anonymous):

What is the perpendicular line for y=.4x+4?

OpenStudy (anonymous):

To be fair, there are infinitely many perpendicular lines to that equation. But anyway, for two lines to be perpendicular, the product of their slopes are -1. Let's call the slope of the new line a. this means that a * 4 = -4, meaning a = -1 So the slope of the new line is -1. The intersection with the Y-axis could be anything, though. They'd still be perpendicular

OpenStudy (anonymous):

Okay, pretty much, there was a question that said a line, y=5/2x, and I had to find a perpendicular line that passed through, (10,0). I got y=.4+4, but I didn't know if that was right. I can'texactly remember how to find slope. I'm homeschooled. Haven't studied that in a year and a half, even though it should be easy.. :c

OpenStudy (phi):

you're pretty close. You want the negative reciprocal of 5/2 that means - 2/5 or -0.4 (not + 0.4) for the slope if the line goes through (10,0) y - 0 = -2/5(x - 10) y = -2/5 x + 4

OpenStudy (anonymous):

Oh.. Thanks. I think I get it now. c: